01 04 2010
Download
Skip this Video
Download Presentation
01/04/2010

Loading in 2 Seconds...

play fullscreen
1 / 14

01/04/2010 - PowerPoint PPT Presentation


  • 85 Views
  • Uploaded on

01/04/2010. 01/04/2010. Start of the studies delayed due to trip of RQD.A12 at 08:43 at the end of pre-cycling  nQPS triggered but no signal observed above threshold.

loader
I am the owner, or an agent authorized to act on behalf of the owner, of the copyrighted work described.
capcha
Download Presentation

PowerPoint Slideshow about ' 01/04/2010' - dennis


An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript
01 04 20101
01/04/2010
  • Start of the studies delayed due to trip of RQD.A12 at 08:43 at the end of pre-cycling  nQPS triggered but no signal observed above threshold.
  • Injection started at ~12:00 for RF capture optimization. Due to the low longitudinal emittance delivered by the SPS the capture voltage could be decreased from 8 to 5 MV at injection
  • Trim to new values implemented in the sequencer
01 04 20102
01/04/2010
  • Preliminary tests of the chromaticity measurement by RF frequency modulation
01 04 20103
01/04/2010
  • Smooth ramp in preparation of squeeze test
  • 15:00 Beam lost at flat-top due to bad manipulation with the sequencer
  • Finally ready to start the squeeze at 18:30
  • No losses during the squeeze (except to over enthusiastic use of AC dipole)
01 04 20104
01/04/2010
  • b* = 9 m
  • Tune and chromaticity OK
01 04 20105
01/04/2010
  • b* = 9 m
  • Beta beating measurement
  • Beam 2
  • Beam 1 (lost 80 % of the beam during AC dipole excitation)
01 04 20106
01/04/2010
  • b* = 7 m
  • Tune and chromaticity OK
01 04 20107
01/04/2010
  • b* = 7 m
  • Beam 2
01 04 20108
01/04/2010
  • b* = 5 m - Tune and chromaticity OK. Coupling increases to ~0.01
01 04 20109
01/04/2010
  • b* = 5 m
  • Beam 2
01 04 201010
01/04/2010
  • Coupling Beam 2
  • b*=9 m b*=5 m
01 04 201011
01/04/2010
  • Important loss during the ramp  beam unstable. Likely chromaticity too low
01 04 201012
01/04/2010
  • Today:
    • 05:00 Ramp (without feedback) - OK
    • 06:37 – lost cryo (oil pump?) – beams dumped
    • access in PZ45 for CRYO intervention
    • around 14:00 – recover from cryo
    • 09:00 Tune feedback / chromaticity measurement during the ramp
    • 14:00 Non linear chromaticity measurement and off-momentum beta beating
    • 17:00-18:00 LHCb polarity change
    • Prepare fill for the night
01 04 201013
01/04/2010
  • Saturday to Monday:
    • 2 fills/day with an end of fill or beginning of fill study (2-3 hours)
  • Beginning/End of fill study list in work
    • Working point in collision
    • Machine protection tests
    • Non-linear chromaticity measurement at flat-top
    • b3 correction checks
ad